# latexOverwriting file

### v4

@@ Appendix A Auxiliary Results
and so the bound (A.4) is valid in general with $L_{\alpha}=\alpha^{K/2}\widetilde{L}_{\alpha^{2}}$.
Step 2: contradiction hypothesis
For simplicity, we use from now on the notations $c,m,\mathit{S}$ in place of $c_{1},m_{1},\mathit{S}_{1}$. It remains to show that there exists for all $\alpha>0$ a constant $\widetilde{L}_{\alpha}$ such that
$\lVert h\rVert_{\infty}\leqslant\widetilde{L}_{\alpha}\lVert h\rVert_{1},\qquad h(y)=cg(y;m,\mathit{S})-g(y;0,\mathit{I}_{K}).$
for all $(c,m,\mathit{S})\in\times^{K}\times\mathcal{S}^{K}_{\alpha}$. Assume for contradiction that there is a sequence of parameters $(c_{n},m_{n},\mathit{S}_{n})_{n\geqslant 0}$ in ×$\times^{K}\times\mathcal{S}^{K}_{\alpha}$ and not containing $(1,0,\mathit{I}_{K})$ and such that
$\frac{\lVert h_{n}\rVert_{\infty}}{\lVert h_{n}\rVert_{1}}\xrightarrow[n\to\infty]{}\infty,\qquad h_{n}(y):=c_{n}g(y;m_{n},\mathit{S}_{n})-g(y;0,\mathit{I}_{K}).$ (A.5)
Since $\lVert h_{n}\rVert_{\infty}\leqslant c_{n}/\sqrt{(2\pi)^{K}\det\mathit{S}_{n}}+1/\sqrt{(2\pi)^{K}}$ and
$\lvert c_{n}-1\rvert=\left\lvert\int_{{}^{K}}h_{n}\,\mathrm{d}y\right\rvert\leqslant\lVert h_{n}\rVert_{1}$
we have
$\frac{\lVert h_{n}\rVert_{\infty}}{\lVert h_{n}\rVert_{1}}\leqslant\frac{c_{n}/\sqrt{\det\mathit{S}_{n}}+1}{\sqrt{(2\pi)^{K}}\lvert c_{n}-1\rvert}\leqslant\frac{c_{n}\sqrt{\alpha^{K}}+1}{\sqrt{(2\pi)^{K}}\lvert c_{n}-1\rvert}.$
The right-hand side can diverge only if $c_{n}\to 1$ in the limit $n\to\infty$, and so this must necessarily hold. If $c_{n}\to 1$ then $\lVert h_{n}\rVert_{\infty}$ is uniformly bounded from above as $n\to\infty$ implying by (A.5) that $\lVert h_{n}\rVert_{1}\to 0$. Now notice that
$h_{n}=\Bigl{(}c_{n}g(y;m_{n},\mathit{S}_{n})-g(y;m_{n},\mathit{S}_{n})\Bigr{)}+\Bigl{(}g(y;m_{n},\mathit{S}_{n})-g(y;0,\mathit{I}_{K})\Bigr{)},$
The $L^{1}(^{K})$ norm of the first term on the right-hand side tends to 0 as $n\to\infty$, so that of the second term must also converge to 0 in $L^{1}(^{K})$. This term may be bounded in norm by using [16, Proposition 2.2], which gives a lower bound on the total variation distance between Gaussians. In our setting, this result implies
$\frac{1}{2}\lVert g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};m_{n},\mathit{S}_{n})-g(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};0,\mathit{I}_{K})\rVert_{1}$ $\geqslant 1-\frac{\det\mathit{S}_{n}^{1/4}\det\mathit{I}_{K}^{1/4}}{\det\left(\frac{\mathit{S}_{n}+\mathit{I}_{K}}{2}\right)^{1/2}}\exp\left(-\frac{1}{8}m_{n}^{\mathsf{T}}\left(\frac{\mathit{S}_{n}+\mathit{I}_{K}}{2}\right)^{-1}m_{n}\right).$ (A.6)
Since the left-hand side of this inequality converges to 0, it necessarily holds that $m_{n}\to 0$ and $\mathit{S}_{n}\to\mathit{I}_{K}$. Indeed, the right-hand side is bounded from below by
$1-\frac{\det\mathit{S}_{n}^{1/4}\det\mathit{I}_{K}^{1/4}}{\det\left(\frac{\mathit{S}_{n}+\mathit{I}_{K}}{2}\right)^{1/2}}=1-\prod_{i=1}^{K}\frac{\lambda_{n,i}^{1/4}}{\left(\frac{\lambda_{n,i}+1}{2}\right)^{1/2}}=1-\prod_{i=1}^{K}\sqrt{\frac{\sqrt{\lambda_{n,i}}}{\frac{\lambda_{n,i}+1}{2}}},$
where $(\lambda_{n,i})_{i=1}^{K}$ are the eigenvalues of $\mathit{S}_{n}$. All the factors in the product are bounded from above by 1 by the arithmetic mean-geometric mean inequality. Since the left-hand side of (A.6) converges to 0, all these factors must converge to 1 in the limit as $n\to\infty$, otherwise the right-hand side of (A.6) would be bounded away from 0 uniformly in $n$. Thus $\lambda_{n,i}\to 1$ for all $i\in\llbracket 1,K\rrbracket$, and so $S_{n}$ converges to $\mathit{I}_{K}$ as $n\to\infty$. This implies that the factor multiplying the exponential in (A.6) converges to 1, so the exponential itself must converge to 1, whence we deduce $m_{n}\to 0$.
Step 3: reaching a contradiction
Since $(c_{n},m_{n},\mathit{S}_{n})$ converges to $(1,0,\mathit{I}_{K})$ as $n\to\infty$, a contradiction can be reached by showing that $\lVert h\rVert_{\infty}/\lVert h\rVert_{1}$ is bounded from above uniformly for any $(c,m,\mathit{S})\neq(1,0,\mathit{I}_{K})$ in a small neighborhood of $(1,0,\mathit{I}_{K})$. To this end, we introduce the following notation for any $(c,m,\mathit{S})\in\times^{K}\times\mathcal{S}^{K}_{\alpha}$:
$\upsilon_{0}=\frac{c}{\sqrt{\det\mathit{S}}}-1,\qquad\upsilon_{1}=m,\qquad\mathit{\Upsilon}_{2}=\frac{\mathit{I}_{K}-\mathit{S}^{-1}}{2},\qquad r=\max\Bigl{\{}\lvert\upsilon_{0}\rvert,\lvert\upsilon_{1}\rvert,\lVert\mathit{\Upsilon}_{2}\rVert\Bigr{\}}.$ (A.7)
Although we omit the arguments $(c,m,\mathit{S})$ for conciseness, these quantities are viewed as functions over ×$\times^{K}\times\mathcal{S}^{K}_{\alpha}$. Then
$h(y)$ $=\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\left((1+\upsilon_{0})\exp\left(-\frac{1}{2}(y-\upsilon_{1})^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})(y-\upsilon_{1})+\frac{y^{\mathsf{T}}y}{2}\right)-1\right).$ (A.8)
The argument of the exponential within the brackets can be rewritten as
$a=a(y;\upsilon_{1},\mathit{\Upsilon}_{2}):=\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y-2\upsilon_{1}^{\mathsf{T}}\mathit{\Upsilon}_{2}y-\frac{1}{2}\upsilon_{1}^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})\upsilon_{1},$
which may be bounded from above as
$\left\lvert a(y;\upsilon_{1},\mathit{\Upsilon}_{2})\right\rvert$ $\leqslant\left(r+r+2r^{2}+\frac{1}{2}r^{2}+r^{3}\right)\bigl{(}1+\lvert y\rvert^{2}\bigr{)},$ (A.9a)
$\left\lvert\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert$ $\leqslant 2r\bigl{(}1+\lvert y\rvert^{2}\bigr{)},$ (A.9b)
$\left\lvert 2\upsilon_{1}^{\mathsf{T}}\mathit{\Upsilon}_{2}y+\frac{1}{2}\upsilon_{1}^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})\upsilon_{1}\right\rvert$ $\leqslant\left(2r^{2}+\frac{1}{2}r^{2}+r^{3}\right)\bigl{(}1+\lvert y\rvert^{2}\bigr{)}$ (A.9c)
Let $\mathcal{A}_{\varepsilon}$ denote the neighborhood $\bigl{\{}(c,m,\mathit{S})\in\times^{K}\times\mathcal{S}^{\alpha}_{K}:r\leqslant\varepsilon\bigr{\}}$. Employing the triangle inequality together with the elementary inequality $\lvert\operatorname{e}^{x}-1-x\rvert\leqslant\frac{x^{2}}{2}\operatorname{e}^{\lvert x\rvert},$ which is valid for any $x\in$ and obtained from a Taylor expansion, we deduce that,
$\forall(c,m,\mathit{S})\in\mathcal{A}_{1},\qquad\forall y\in^{K},\qquad$
$\qquad\left\lvert\operatorname{e}^{a}-1-\upsilon_{1}^{\mathsf{T}}y-y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert$ $\leqslant\left\lvert 2\upsilon_{1}^{\mathsf{T}}\mathit{\Upsilon}_{2}y+\frac{1}{2}\upsilon_{1}^{\mathsf{T}}(\mathit{I}_{K}-2\mathit{\Upsilon}_{2})\upsilon_{1}\right\rvert+\frac{1}{2}a^{2}\operatorname{e}^{\lvert a\rvert}$
$\leqslant 4r^{2}\left(1+\lvert y\rvert^{2}\right)+\frac{1}{2}(6r)^{2}\bigl{(}1+\lvert y\rvert^{2}\bigr{)}^{2}\exp\biggl{(}6r\left(1+\lvert y\rvert^{2}\right)\biggr{)}$
$\leqslant 22r^{2}\left(1+\lvert y\rvert^{2}\right)^{2}\exp\biggl{(}6r\left(1+\lvert y\rvert^{2}\right)\biggr{)}.$
We have used the fact that, since $(c,m,\mathit{S})\in\mathcal{A}_{1}$, we have $r\leqslant 1.$ From this equation and (A.8) and the expression
$h(y)-\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\left(\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right)$
$\qquad=\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\Bigl{(}\operatorname{e}^{a}-1-\upsilon_{1}^{\mathsf{T}}y-y^{\mathsf{T}}\mathit{\Upsilon}_{2}y+\upsilon_{0}(\operatorname{e}^{a}-1)\bigr{)}\Bigr{)},$
@@
we deduce that there exists $\varepsilon_{1}\in(0,1)$ sufficiently small such that for all $(c,m,\mathit{S})\in\mathcal{A}_{\varepsilon_{1}}$, the following inequality holds:
$\forall y\in^{K},\qquad\left\lvert h(y)-\frac{\exp\left(-\frac{y^{\mathsf{T}}y}{2}\right)}{\sqrt{(2\pi)^{K}}}\left(\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right)\right\rvert\leqslant Cr^{2}\exp\left(-\frac{y^{\mathsf{T}}y}{4}\right).$ (A.10)
Therefore, it holds that $\lVert h\rVert_{\infty}$ is bounded from above by $Cr$ uniformly for $(c,m,\mathit{S})$ in $\mathcal{A}_{\varepsilon_{1}}$. In order to reach a contradiction and conclude the proof, it is therefore sufficient to prove that $\lVert h\rVert_{1}$ is bounded from below by $Cr$ for $(c,m,\mathit{S})$ in $\mathcal{A}_{\varepsilon_{2}}$ for some $\varepsilon_{2}\in(0,1)$. It follows from (A.10) that
$\lVert h\rVert_{1}=\int_{{}^{K}}\left\lvert\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert\,g(y;0,\mathit{I}_{K})\,\mathrm{d}y+\mathcal{O}(r^{2}).$
Let $\mathcal{I}$ denote the integral on the right-hand side. It holds that
$\frac{\mathcal{I}}{r}=\int_{{}^{K}}\left\lvert\left(\frac{\upsilon_{0}}{r}\right)+\left(\frac{\upsilon_{1}}{r}\right)^{\mathsf{T}}y+y^{\mathsf{T}}\left(\frac{\mathit{\Upsilon}_{2}}{r}\right)y\right\rvert\,g(y;0,\mathit{I}_{K})\,\mathrm{d}y.$
By definition of $r$, all of the coefficients in parentheses are bounded from above by $1$, in absolute value, vector norm and matrix norm respectively; furthermore one of the terms is necessarily equal to $1$ by one of these measures of size. The conclusion then follows from the lower bound (A.11), which is proved in the next step.
Step 4: proof of the auxiliary inequality
In this step, we prove that there exists a constant $C>0$ such that the following inequality holds for all $(\upsilon_{0},\upsilon_{1},\mathit{\Upsilon}_{2})\in\times^{K}\times^{K\times K}$ satisfying $\max\bigl{\{}\lvert\upsilon_{0}\rvert,\lvert\upsilon_{1}\rvert,\lVert\mathit{\Upsilon}_{2}\rVert\bigr{\}}=1$:
@@
$\int_{{}^{K}}\left\lvert\upsilon_{0}+\upsilon_{1}^{\mathsf{T}}y+y^{\mathsf{T}}\mathit{\Upsilon}_{2}y\right\rvert\,g(y;0,\mathit{I}_{K})\mathrm{d}y\geqslant C.$ (A.11)
To this end, note that set of parameters $(\upsilon_{0},\upsilon_{1},\mathit{\Upsilon}_{2})$ satisfying $\max\bigl{\{}\lvert\upsilon_{0}\rvert,\lvert\upsilon_{1}\rvert,\lVert\mathit{\Upsilon}_{2}\rVert\bigr{\}}=1$ is a compact subset, say $\mathcal{K}$, of ×$\times^{K}\times^{K\times K}$. In addition, the left-hand side of (A.11) is continuous with respect to the parameters. Therefore, the infimum of the integral over $\mathcal{K}$ is attained, and since the integral is strictly positive for any tuple of parameters in $\mathcal{K}$, the infimum is necessarily strictly positive. ∎
@@
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context, depends only on $\kappa_{\Psi},\kappa_{h},\ell_{h},\mathit{\Sigma},\mathit{\Gamma}$, and may change from line to line. We begin by proving that for all $(u_{1},u_{2},v)\in^{d}\times^{d}\times^{d}$, it holds that
@@
$\hskip 142.26378pt\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.13)
Since the operator $\Psi$ is bounded by Assumption H, there is $C$ such that
$\forall(u,v)\in^{d}\times^{d},\qquad\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\leqslant C\exp\left(-\frac{1}{3}\lvert u\rvert_{\mathit{\Sigma}}^{2}\right),$ (A.14)
because the ratio between the exponential functions on the left and right-hand side converges to 00 in the limit as $\lvert u\rvert\to\infty$. By (A.14) it is clear that there is $C>0$ such (A.13) is satisfied whenever $\lvert u_{2}-u_{1}\rvert\geqslant 1$. We now prove that there is $C>0$ such that (A.13) is satisfied when $\lvert u_{2}-u_{1}\rvert\leqslant 1$. It follows from (A.14) and the elementary inequality
$\forall(a,b)\in\times,\qquad\left\lvert\operatorname{e}^{-a}-\operatorname{e}^{-b}\right\rvert\leqslant\lvert b-a\rvert\operatorname{e}^{-\min\{a,b\}}$ (A.15)
that, for any $(u_{1},u_{2},v)\in^{d}\times^{d}\times^{d}$,
$\qquad\leqslant\frac{C}{2}\left\lvert\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}-\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right\rvert\exp\left(-\frac{1}{3}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\left\lvert u_{2}\right\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.16)
Using that $a^{2}-b^{2}=(a-b)(a+b)$ and then the triangle inequality, we obtain
$\left\lvert\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}-\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right\rvert\leqslant\lvert u_{1}-u_{2}\rvert_{\mathit{\Sigma}}\Bigl{(}\lvert u_{1}\rvert_{\mathit{\Sigma}}+\lvert u_{2}\rvert_{\mathit{\Sigma}}+2\lvert\Psi(v)\rvert_{\mathit{\Sigma}}\Bigr{)}.$ (A.17)
The norm $\lvert\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\rvert_{\mathit{\Sigma}}$ is equivalent to the Euclidean norm, and so the right-hand side of (A.16) is bounded from above when $\lvert u_{2}-u_{1}\rvert\leqslant 1$ by
$C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$
Note that the factor in the exponential is now $1/4$ instead of $1/3$, because we used part of the exponential decay in (A.16) in order to control the growth of the term $\lvert u_{1}\rvert_{\mathit{\Sigma}}+\lvert u_{2}\rvert_{\mathit{\Sigma}}$, which appears on the right-hand side of (A.17). It then follows from (A.13) and the definition (1.5) of $\mathsf{P}$ that the density of $\mathsf{P}\mu$ satisfies the following inequality for all $(u_{1},u_{2})\in^{d}\times^{d}$:
@@
$\bigl{\lvert}\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\bigr{\rvert}\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.18)
The above reasoning is now repeated almost verbatim in order to prove that the following inequality holds for all $(u_{1},u_{2},y)\in^{d}\times^{K}\times^{K}$:
$\Bigl{\lvert}\mathcal{N}\bigl{(}h(u_{1}),\mathit{\Gamma}\bigr{)}(y)-\mathcal{N}\bigl{(}h(u_{2}),\mathit{\Gamma}\bigr{)}(y)\Bigr{\rvert}\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right).$ (A.19)
Indeed, since $h$ is bounded by Assumption H, it holds that
$\forall(u,y)\in^{d}\times^{K},\qquad\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)$ $\leqslant C\exp\left(-\frac{1}{3}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right),$ (A.20)
and so by (A.15) we have

### v6

@@ Appendix A Auxiliary Results
and so the bound (A.6) is valid in general with $\widetilde{L}_{\alpha}=\alpha^{\frac{K}{2}}\widehat{L}_{\alpha^{2}}$.
Step 2. Proof of the simplified statement
@@
It remains to show that there exists for all $\alpha>0$ a constant $\widehat{L}_{\alpha}$ such that the following inequality holds for all $(m,\mathit{S})\in^{K}\times\mathcal{S}^{K}_{\alpha}$:
$\lVert\mathfrak{h}\rVert_{\infty}\leqslant\widehat{L}_{\alpha}\lVert\mathfrak{h}\rVert_{1},\qquad\mathfrak{h}(y)=g(y;m,\mathit{S})-g(y;0,\mathit{I}_{K}).$ (A.7)
To this end, fix $\alpha>0$, fix $\varepsilon\in(0,1)$ and assume first that $\lVert\mathit{S}-\mathit{I}_{K}\rVert\geqslant\varepsilon$. By the lower bound on the total variation distance between Gaussians in [2018arXiv181008693D, Proposition 2.2], we have that
@@
$\frac{1}{2}\lVert\mathfrak{h}\rVert_{1}$ $\geqslant 1-\frac{\det\mathit{S}^{1/4}\det\mathit{I}_{K}^{1/4}}{\det\left(\frac{\mathit{S}+\mathit{I}_{K}}{2}\right)^{1/2}}\exp\left(-\frac{1}{8}m^{\mathsf{T}}\left(\frac{\mathit{S}+\mathit{I}_{K}}{2}\right)^{-1}m\right).$
In particular, it holds that
$\frac{1}{2}\lVert\mathfrak{h}\rVert_{1}$ $\geqslant 1-\frac{\det\mathit{S}^{1/4}\det\mathit{I}_{K}^{1/4}}{\det\left(\frac{\mathit{S}+\mathit{I}_{K}}{2}\right)^{1/2}}=1-\prod_{i=1}^{K}\sqrt{\frac{\sqrt{\lambda_{i}}}{\frac{\lambda_{i}+1}{2}}}\geqslant 1-\sqrt{\frac{2\sqrt{1+\varepsilon}}{2+\varepsilon}},$ (A.8)
where $(\lambda_{i})_{i=1}^{K}$ are the eigenvalues of $\mathit{S}$. In the last inequality, we used that, since all the terms in the product are bounded from above by 1 by the arithmetic mean-geometric mean inequality, and since at least one eigenvalue is not in the interval $(1-\varepsilon,1+\varepsilon)$ given that $\lVert\mathit{S}-\mathit{I}_{K}\rVert\geqslant\varepsilon$, it holds that
$\prod_{i=1}^{K}\sqrt{\frac{\sqrt{\lambda_{i}}}{\frac{\lambda_{i}+1}{2}}}\leqslant\max_{|\lambda-1|\geqslant\varepsilon}\sqrt{\frac{\sqrt{\lambda}}{\frac{\lambda+1}{2}}}=\sqrt{\frac{2\sqrt{1+\varepsilon}}{2+\varepsilon}}.$
On the other hand, since $S\succcurlyeq\frac{1}{\alpha}I_{K}$, it holds that $\det S\geqslant\frac{1}{\alpha^{K}}$, and so
$\lVert\mathfrak{h}\rVert_{\infty}\leqslant 2\left(\frac{\alpha}{2\pi}\right)^{\frac{K}{2}}.$ (A.9)
Combining (A.8) and (A.9) gives that
$\lVert\mathfrak{h}\rVert_{1}\geqslant C_{1}\lVert\mathfrak{h}\rVert_{\infty},\qquad C_{1}:=\left(1-\sqrt{\frac{2\sqrt{1+\varepsilon}}{2+\varepsilon}}\right)\left(\frac{2\pi}{\alpha}\right)^{\frac{K}{2}}.$
Consider now the case where $\lVert\mathit{S}-\mathit{I}_{K}\rVert\leqslant\varepsilon$. Since $S\mapsto\sqrt{\det S}$ is Lipschitz continuous over the set of symmetric positive definite matrices $S$ such that $\lVert\mathit{S}-\mathit{I}_{K}\rVert\leqslant\varepsilon$, with a Lipschitz constant we denote by $c_{\varepsilon}$, it holds by Lemma 13 that
$\lVert\mathfrak{h}\rVert_{\infty}$ $\leqslant(2\pi)^{-\frac{K}{2}}\left(\frac{\alpha^{\frac{1+K}{2}}}{\sqrt{\operatorname{e}}}\lvert m\rvert+\frac{\alpha^{1+\frac{K}{2}}}{\operatorname{e}}\lVert\mathit{S}-\mathit{I}_{K}\rVert+\alpha^{K}\left\lvert\sqrt{\det S}-\sqrt{\det\mathit{I}_{K}}\right\rvert\right)$
$\leqslant(2\pi)^{-\frac{K}{2}}\left(\frac{\alpha^{\frac{1+K}{2}}}{\sqrt{\operatorname{e}}}\lvert m\rvert+\left(\frac{\alpha^{1+\frac{K}{2}}}{\operatorname{e}}+c_{\varepsilon}\alpha^{K}\right)\lVert\mathit{S}-\mathit{I}_{K}\rVert\right).$
In view of (A.9), this implies that there exists a constant $C$ depending only on $(\varepsilon,\alpha,K)$ such that
$\lVert\mathfrak{h}\rVert_{\infty}\leqslant C\Bigl{(}\min\bigl{\{}\lvert m\rvert,1\bigr{\}}+\lVert\mathit{S}-\mathit{I}_{K}\rVert\Bigr{)}.$ (A.10)
On the other hand, since the characteristic function of $\mathcal{N}(m,\mathit{S})$ is given by $u\mapsto\operatorname{e}^{{\rm i}m^{\mathsf{T}}u-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}$, it holds by definition of the characteristic function that
$\forall u\in^{K},\qquad\operatorname{e}^{{\rm i}m^{\mathsf{T}}u-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}-\operatorname{e}^{-\frac{\lvert u\rvert^{2}}{2}}=\int_{{}^{K}}\operatorname{e}^{{\rm i}u^{\mathsf{T}}x}\mathfrak{h}(x)\,\mathrm{d}x.$
Therefore, it holds that
$\lVert\mathfrak{h}\rVert_{1}$ $\geqslant\sup_{|u|\leqslant 1}\left\lvert\operatorname{e}^{{\rm i}m^{\mathsf{T}}u-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}-\operatorname{e}^{-\frac{\lvert u\rvert^{2}}{2}}\right\rvert=\sup_{|u|\leqslant 1}\left\lvert\operatorname{e}^{-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}-\operatorname{e}^{-{\rm i}m^{\mathsf{T}}u-\frac{\lvert u\rvert^{2}}{2}}\right\rvert.$
It is clear from elementary geometry in the complex plane that
$\forall u\in^{K},\qquad\left\lvert\operatorname{e}^{-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}-\operatorname{e}^{-{\rm i}m^{\mathsf{T}}u-\frac{\lvert u\rvert^{2}}{2}}\right\rvert\geqslant\max\left\{\left\lvert\sin(m^{\mathsf{T}}u)\right\rvert\operatorname{e}^{-\frac{\lvert u\rvert^{2}}{2}},\left\lvert\operatorname{e}^{-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}-\operatorname{e}^{-\frac{\lvert u\rvert^{2}}{2}}\right\rvert\right\},$
and so we have
$\lVert\mathfrak{h}\rVert_{1}$ $\geqslant\max\left\{\sup_{|u|\leqslant 1}\left\lvert\sin(m^{\mathsf{T}}u)\right\rvert\operatorname{e}^{-\frac{\lvert u\rvert^{2}}{2}},\sup_{|u|\leqslant 1}\left\lvert\operatorname{e}^{-\frac{1}{2}u^{\mathsf{T}}\mathit{S}u}-\operatorname{e}^{-\frac{\lvert u\rvert^{2}}{2}}\right\rvert\right\}$
$\geqslant\max\left\{\operatorname{e}^{-\frac{1}{2}}\sup_{|u|\leqslant 1}\left\lvert\sin(m^{\mathsf{T}}u)\right\rvert,\frac{\operatorname{e}^{-1}}{2}\sup_{|u|\leqslant 1}\left\lvert u^{\mathsf{T}}(\mathit{S}-\mathit{I}_{K})u\right\rvert\right\}$
$=\max\left\{\operatorname{e}^{-\frac{1}{2}}\sup_{|u|\leqslant 1}\left\lvert\sin(m^{\mathsf{T}}u)\right\rvert,\frac{\operatorname{e}^{-1}}{2}\lVert\mathit{S}-\mathit{I}_{K}\rVert\right\}.$
In the second inequality, we used that $u^{\mathsf{T}}Su\leqslant 1+\varepsilon\leqslant 2$ for all $|u|\leqslant 1$, together with the elementary inequality $\left\lvert\operatorname{e}^{a}-\operatorname{e}^{b}\right\rvert\geqslant\operatorname{e}^{\min\{a,b\}}|a-b|$. To conclude, considering the particular value
$u=\frac{m}{|m|\max\left\{1,\frac{4|m|}{\pi}\right\}},$
we obtain that
$\sup_{|u|\leqslant 1}\left\lvert\sin(m^{\mathsf{T}}u)\right\rvert\geqslant\sin\left(\min\left\{|m|,\frac{\pi}{4}\right\}\right)=\int_{0}^{\min\left\{|m|,\frac{\pi}{4}\right\}}\cos(t)\,\mathrm{d}t\geqslant\cos\left(\frac{\pi}{4}\right)\min\left\{|m|,\frac{\pi}{4}\right\}.$
Thus, using that $\max\{A,B\}\geqslant\frac{A}{2}+\frac{B}{2}$, we obtain
$\lVert\mathfrak{h}\rVert_{1}$ $\geqslant\max\left\{\operatorname{e}^{-\frac{1}{2}}\cos\left(\frac{\pi}{4}\right)\min\left\{|m|,\frac{\pi}{4}\right\},\frac{\operatorname{e}^{-1}}{2}\lVert\mathit{S}-\mathit{I}_{K}\rVert\right\}$
$\geqslant\frac{1}{2}\min\Bigl{\{}\operatorname{e}^{-\frac{1}{2}}\cos\left(\frac{\pi}{4}\right)\frac{\pi}{4},\frac{\operatorname{e}^{-1}}{2}\Bigr{\}}\Bigl{(}\min\left\{\lvert m\rvert,1\right\}+\lVert\mathit{S}-\mathit{I}_{K}\rVert\Bigr{)}$ (A.11)
Combining (A.11) with (A.10) leads to $\lVert\mathfrak{h}\rVert_{1}\geqslant C_{2}\lVert\mathfrak{h}\rVert_{\infty}$, which concludes the proof of the case $\lVert\mathit{S}-\mathit{I}_{K}\rVert\leqslant\varepsilon$. Consequently, the statement (A.7) holds in general with constant $\widehat{L}_{\alpha}=\min\{C_{1},C_{2}\}^{-1}$. ∎
@@
Throughout this proof, $C$ denotes a constant whose value is irrelevant in the context, depends only on $\kappa_{\Psi},\kappa_{h},\ell_{h},\mathit{\Sigma},\mathit{\Gamma}$, and may change from line to line. Since the function $g(x):=\operatorname{e}^{-x^{2}}$ has derivative $-2x\operatorname{e}^{-x^{2}}$ and since $\lvert x\operatorname{e}^{-x^{2}}\rvert\leqslant\operatorname{e}^{-\frac{2x^{2}}{3}}$ for all $x\in$, it holds for all $(a,b)\in^{2}$ that there is $\xi$ between $|a|$ and $|b|$ such that
$\left\lvert\operatorname{e}^{-a^{2}}-\operatorname{e}^{-b^{2}}\right\rvert=\lvert b-a\rvert\,\lvert g^{\prime}(\xi)\rvert\leqslant 2\lvert b-a\rvert\left(\operatorname{e}^{-\frac{2a^{2}}{3}}+\operatorname{e}^{-\frac{2b^{2}}{3}}\right).$ (A.13)
Using this inequality with $a^{2}=\frac{1}{2}\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}$ and $b^{2}=\frac{1}{2}\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}$, and then using the triangle inequality, we deduce that, for all $(u_{1},u_{2},v)\in^{d}\times^{d}\times^{d}$,
@@
$\hskip 56.9055pt\leqslant C\lvert u_{2}-u_{1}\rvert_{\mathit{\Sigma}}\left(\exp\left(-\frac{1}{3}\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)+\exp\left(-\frac{1}{3}\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\right).$
Integrating out the $v$ variable with respect to $\mu$ and using the equivalence of norms, we obtain that
@@
$\lvert\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\rvert$ $\leqslant C\lvert u_{1}-u_{2}\rvert\int_{{}^{d}}\exp\left(-\frac{1}{3}\left\lvert u_{1}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\,\mu(\mathrm{d}v)$